### v1

@@ 4.4.1 Quantitative analysis
This section assesses the segmentation performance of MSA-UNet3+ on the coronary DSA dataset. The quantitative results in Table 2 demonstrate the superior performance of the model, particularly in capturing fine vascular structures compared to existing methods. Evaluation of our proposed MSA-UNet3+ on the coronary DSA dataset (Table 2) reveals competitive performance across key metrics. Although MCDAU_Net attains the highest Recall ($87.71\pm 0.0276$), MSA-UNet3+ achieves superior performance in several critical aspects: the second-highest Recall ($86.78\pm 0.0080$), the best F1 score ($87.78\pm 0.0036$) and Dice coefficient ($87.73\pm 0.0036$) among all models, and exceptional boundary precision with the lowest ASD ($0.76\pm 0.0414$) and ACD ($0.74\pm 0.0367$). While CMU_Net (Recall: $85.14\pm 0.0209$) and MCDAU_Net demonstrate competitive results, they underperform MSA-UNet3+ in overall F1, Dice, and boundary accuracy Similarly, FR_UNet and CA_Net show competent but consistently inferior performance relative to our proposed model. These results establish MSA-UNet3+ as a state-of-the-art solution for coronary DSA segmentation, combining a balanced metric performance with clinically crucial boundary precision.
@@
Fig. 8 illustrates the trade-off between model efficiency (measured by the numbers of parameters) and segmentation accuracy (quantified by DSC and ACD). The proposed MSA-UNet3+ achieves an optimal balance, attaining a DSC of 87.78% with only 7.54 million parameters—significantly fewer than comparably accurate models such as CA_Net, CMU_Net, and MMDC_Net. While PMFS_Net demonstrates greater efficiency (0.33M parameters), its lower DSC (84.49%) highlights the accuracy-efficiency compromise. Visual encoding in the figure enhances interpretation: the point size scales with model size, while the color intensity reflects ASD, with darker colors indicating lower ASD (better surface distance accuracy). MSA-UNet3+ excels on both axes, combining competitive ASD (0.76) with compact architecture. These results position MSA-UNet3+ as an ideal solution for practical applications where computational resources and segmentation quality are critical.
@@
Further validating the model’s robustness, the performance distribution analysis in Fig. 9 demonstrates that our approach (“ours”) achieves both a high median Dice coefficient and a relatively tight interquartile range, suggesting reliable segmentation performance. The visualization’s annotated markers (orange circles for medians, green triangles for means) reveal an important trade-off: while models like CMU_Net and MMDC_Net show higher median Dice coefficients but with wider interquartile ranges, indicating less consistency in performance.
In contrast, more stable models (MGA_Net, MCDAU_Net) achieve narrower ranges but with reduced segmentation quality. Together with the parameter efficiency analysis, these results position our solution as clinically optimal, balancing performance, consistency, and computational practicality for clinical implementation.
@@ 4.4.2 Qualitative analysis
Detailed analysis reveals that while BCU_Net and CMU_NeXt produce substantial false positive detections in complex anatomical backgrounds, DATrans_Unet and Isunetv1 exhibit pronounced false negative rates, failing to identify fine vascular details. PMFS_Net demonstrates intermediate performance with variable accuracy across different vascular structures.
The close visual correspondence between our segmentation results and the ground truth annotations highlights the model’s potential for clinical implementation in coronary artery analysis, where precise vessel delineation is critical for accurate diagnosis and treatment planning.
Table 3: Ablation study of the proposed MSA-UNet3+ framework, presenting performance metrics for eight configurations to assess the contributions of the MSD-Bottleneck, CAFM, and SPCL modules. Symbols denote: (✓) presence and (✗) absence of each component in the ablation study.
@@
Idx PSCE CAFM MSD Recall $\uparrow$ F1 $\uparrow$ Dice $\uparrow$ ASD $\downarrow$ ACD $\downarrow$
@@ 4.5.1 Effect of the Supervised Prototypical Contrastive Loss (SPCL)
The SPCL represents a specialized loss function engineered to improve feature discrimination by promoting more distinct class representations within the learned feature space. When incorporated into the baseline model, SPCL generates consistent improvements across all evaluation metrics: Recall exhibits an increase from 85.212% to 85.305%, while both the F1-score and Dice coefficient show measurable gains, improving from 87.233% and 87.181% to 87.418% and 87.367%, respectively. Most notably, SPCL produces substantial reductions in segmentation error metri- cs,with the Average Surface Distance (ASD) decreasing from 0.8914 to 0.8517 and the Average Contour Distance (ACD) improving from 0.8639 to 0.8263. These results demonstrate SPCL’s ability to sharpen class separation, particularly at challenging boundaries, by refining the model’s feature representations. The complete potential of SPCL becomes particularly apparent when integrated with complementary architectural components. For example, when combined with the Multi-Scale Dilated Bottleneck (baseline+MSD+SPCL configuration), SPCL enables significant performance enhancements. The F1-score and Dice coefficient rise to 87.726% and 87.680% respectively, while the ASD and ACD metrics demonstrate further improvement, reaching 0.7689 and 0.7509. These results represent a marked advancement over the baseline performance. This combination highlights SPCL’s ability to complement multi-scale feature extraction, enabling the model to distinguish between classes with remarkable precision across varying scales. SPCL contributes significantly by improving feature discrimination and enhancing multi-scale capabilities.
4.5.2 Impact of the MSD-Bottleneck
@@
The MSD-Bottleneck module employs dilated convolutions with varying dilation rates to capture multi-scale features, thereby enabling robust handling of objects with diverse sizes and morphological characteristics, a critical capability for coronary DSA image segmentation. When integrated into the baseline architecture, the MSD-Bottleneck produces substantial performance gains, increasing Recall from 85.212% to 86.530% while simultaneously reducing both the Average Surface Distance (ASD) from 0.8914 to 0.7989 and the Average Contour Distance (ACD) from 0.8639 to 0.8136. These quantitative improvements underscore the fundamental importance of multi-scale feature extraction for accurate detection and segmentation of vascular structures across different spatial scales. The module’s effectiveness is further amplified when combined with the Supervised Prototypical Contrastive Loss (SPCL), as evidenced by additional reductions in segmentation error metrics to 0.7689 (ASD) and 0.7509 (ACD).
@@
This complementary interaction between MSD-Bottleneck and SPCL arises from their complementary mechanisms: while MSD-Bottleneck provides comprehensive multi-scale feature extraction through its dilated convolution architecture, SPCL concurrently optimizes the discriminative quality of these features. The combined operation of these components yields superior segmentation performance compared to their individual contributions, demonstrating that multi-scale feature capture and feature space refinement operate in a mutually reinforcing manner.
@@ 4.5.3 Role of the Contextual Attention Fusion Module (CAFM)
The Contextual Attention Fusion Module (CAFM) enhances segmentation performance by selectively integrating contextual information from diverse image regions while suppressing irrelevant features. This attention mechanism substantially improves the model’s global context comprehension, a critical factor for precise coronary vessel segmentation. When implemented with the baseline architecture (baseline+CAFM), the module demonstrates measurable performance improvements: Recall increases from 85.212% to 85.990%, accompanied by rises in both F1-score (87.233% to 87.593%) and Dice coefficient (87.181% to 87.547%). Furthermore, CAFM reduces segmentation errors substantially, as evidenced by decreases in Average Surface Distance (ASD) from 0.8914 to 0.8003 and Average Contour Distance (ACD) from 0.8639 to 0.7820, confirming enhanced boundary alignment precision. However, the module’s incremental benefits become less substantial when combined with other advanced components. In configurations combining CAFM with either the Supervised Prototypical Contrastive Loss (baseline+SPCL+CAFM) or the Multi-Scale Dilated Bottleneck (baseline+CAFM+MSD), performance gains are more modest, showing only slight improvements in F1-scores and minima- l reductions in ASD and ACD metrics. This observation suggests that while CAFM independently provides significant contextual understanding enhancements, its additive value diminishes when integrated with other sophisticated modules, as their respective contributions appear to operate through complementary rather than amplifying mechanisms with respect to CAFM’s functionality.
@@ 4.5.4 Combined Contributions
The complete MSA-UNet3+ architecture, integrating all three key components (MSD-Bottleneck, CAFM, and SPCL), demonstrates optimal performance across all evaluation metrics. The unified model achieves a Recall of 86.775%, along with an F1-score of 87.776% and Dice coefficient of 87.733%, while simultaneously attaining the most favorable boundary alignment metrics with ASD and ACD values of 0.7587 and 0.7411 respectively. These results illustrate the effective collaborative operation of the constituent modules: the MSD-Bottleneck’s multi-scale feature extraction capability, CAFM’s contextual information integration, and SPCL’s enhanced feature discrimination collectively establish a comprehensive and precise segmentation framework. Each component addresses specific challenges in coronary vessel segmentation, and their combined operation enables superior performance in handling complex anatomical variations. Through systematic ablation analysis, we confirm that every module (MSD-Bottleneck, CAFM, and SPCL) makes substantial and distinct contributions to the overall performance of MSA-UNet3+. The complete integration of these components yields segmentation results that surpass current state-of-the-art approaches, positioning MSA-UNet3+ as a particularly effective solution for the demanding requirements of coronary DSA image analysis. The framework’s success stems from its balanced combination of multi-scale processing, contextual awareness, and discriminative feature learning, which together address the principal challenges in vascular segmentation tasks.
@@ 5 Conclusion
To address the challenges of inadequate contextual information and loss of microvascular features in coronary DSA image segmentation, we propose a novel Multi-Scale Attention modified UNet3+ (MSA-UNet3+) framework. This framework integrates a Multi-Scale Dilated Bottleneck (MSD-Bottleneck) and a Contextual Attention Fusion Module (CAFM) to enhance multi-scale feature extraction and contextual understanding. Additionally, we introduce a novel Supervised Prototypical Contrastive Loss (SPCL) to mitigate class imbalance and high intra-class variance by focusing on hard-to-classify background samples. The MSA-UNet3+ framework effectively captures both fine-grained details and broader structural information, enabling precise segmentation of coronary arteries in DSA images. Our approach has been evaluated through qualitative and quantitative analyses on a coronary DSA dataset, demonstratin- g significant improvements in segmentation performance, particularly for low-contrast and small vessels. The proposed model achieves state-of-the-art performance, with superior Dice coefficients, F1-scores, recall, and reduced boundary errors compared to existing methods.
@@
While the proposed framework demonstrates promising segmentation performance, several limitations warrant consideration. The model’s effectiveness remains constrained when processing DSA images containing severe noise contamination, significant motion artifacts, or extremely low contrast conditions where vascular structures become fundamentally indistinct. These challenging scenarios currently lead to suboptimal segmentation outcomes due to the inherent limitations of angiographic imaging. Future research directions will focus on three key improvements: First, we will investigate multimodal fusion approaches incorporating complementary imaging modalities such as intravascular ultrasound (IVUS) and optical coherence tomography (OCT) to enhance segmentation reliability in problematic cases. Second, we plan to develop optimized lightweight variants of the architecture suitable for deployment on medical imaging devices, thereby facilitating real-time clinical decision-making during interventional procedures. Third, although the current implementation specifically targets coronary DSA analysis, the underlying methodology shows considerable potential for adaptation to other medical segmentation tasks characterized by similar challenges, including class imbalance and complex anatomical backgrounds.
6 Declaration of Competing Interest
The authors declare they have no conflicts of interest.
7 Acknowledgements
@@
The key project of the Shenzhen Science and Technology Innovation Bureau (No. JSGG20220831110400001, No. CJGJZD20230724093303007, KJZD20240903101259-001), is funded in part by Shenzhen Medical Research Fund (No. D2404001), is funded in part by Shenzhen Engineering Laboratory for the Diagnosis & Treatment Key Technologies of Interventional Surgical Robots (XMHT20220104009), and the Key Laboratory of Biomedical Imaging Science and System, CAS.
8 Data availability

### v3

@@ 4.4.1 Quantitative analysis
This section assesses the segmentation performance of MSA-UNet3+ with SPCL loss on the coronary DSA dataset. The quantitative results in Table 2 demonstrate the superior performance of the model, particularly in capturing fine vascular structures compared to existing methods. Evaluation of our proposed MSA-UNet3+ on the coronary DSA dataset (Table 2) reveals competitive performance across key metrics. Although MCDAU_Net attains the highest recall ($87.71\pm 0.0276$), MSA-UNet3+ achieves superior performance in several critical aspects: the second-highest recall ($86.78\pm 0.0080$), the best F1 score ($87.78\pm 0.0036$) and Dice coefficient ($87.73\pm 0.0036$) among all models, and exceptional boundary precision with the lowest ASD ($0.76\pm 0.0414$) and ACD ($0.74\pm 0.0367$). While CMU_Net (recall: $85.14\pm 0.0209$) and MCDAU_Net demonstrate competitive results, they underperform MSA-UNet3+ in overall F1, Dice, and boundary accuracy. Similarly, FR_UNet and CA_Net show competent but consistently inferior performance relative to our proposed model. These results establish MSA-UNet3+ as a state-of-the-art solution for coronary DSA segmentation, combining a balanced metric performance with clinically crucial boundary precision.
@@
Fig. 9 illustrates the trade-off between model efficiency (measured by the number of parameters) and segmentation accuracy (quantified by DSC and ASD). The proposed MSA-UNet3+ achieves an optimal balance, attaining a DSC of 87.78% with only 7.54 million parameters—significantly fewer than comparably accurate models such as CA_Net, CMU_Net, and MMDC_Net. While PMFS_Net demonstrates greater efficiency (0.33M parameters), its lower DSC (84.49%) highlights the accuracy-efficiency compromise. Visual encoding in the figure enhances interpretation: the point size scales with model size, while the color intensity reflects ASD, with darker colors indicating lower ASD (better surface distance accuracy). MSA-UNet3+ excels on both axes, combining competitive ASD (0.76) with compact architecture. These results position MSA-UNet3+ as an ideal solution for practical applications where computational resources and segmentation quality are critical.
@@
Further validating the model’s robustness, the performance distribution analysis in Fig. 10 demonstrates that our approach (“ours”) achieves both a high median Dice coefficient and a relatively tight interquartile range, suggesting reliable segmentation performance. The visualization’s annotated markers (orange circles for medians, green triangles for means) reveal an important trade-off: while models like CMU_Net and MMDC_Net show higher median Dice coefficients, they have wider interquartile ranges, indicating less consistency in performance. In contrast, more stable models (MGA_Net, MCDAU_Net) achieve narrower ranges but with reduced segmentation quality. Together with the parameter efficiency analysis, these results position our solution as clinically optimal, balancing performance, consistency, and computational practicality for clinical implementation.
@@ 4.4.2 Qualitative analysis
Detailed analysis reveals that while BCU_Net and CMU_NeXt produce substantial false positive detections in complex anatomical backgrounds, DATrans_Unet and Isunetv1 exhibit pronounced false negative rates, failing to identify fine vascular details. PMFS_Net demonstrates intermediate performance with variable accuracy across different vascular structures. The close visual correspondence between our segmentation results and the ground truth annotations highlights the model’s potential for clinical implementation in coronary artery analysis, where precise vessel delineation is critical for accurate diagnosis and treatment planning.
Table 3: Ablation study of the proposed MSA-UNet3+ framework, presenting performance metrics for eight configurations to assess the contributions of the MSD-Bottleneck, CAFM, and SPCL modules. Symbols denote (✓) presence and (✗) absence of each component in the ablation study.
@@
Idx SPCL CAFM MSD Recall $\uparrow$ F1 $\uparrow$ Dice $\uparrow$ ASD $\downarrow$ ACD $\downarrow$
@@ 4.5.1 Effect of the supervised prototypical contrastive loss (SPCL)
The SPCL represents a specialized loss function engineered to improve feature discrimination by promoting more distinct class representations within the learned feature space. When incorporated into the baseline model, SPCL generates consistent improvements across all evaluation metrics: Recall exhibits an increase from 85.212% to 85.305%, while both the F1-score and Dice coefficient show measurable gains, improving from 87.233% and 87.181% to 87.418% and 87.367%, respectively. Most notably, SPCL produces substantial reductions in segmentation error met- rics, with the Average Surface Distance (ASD) decreasing from 0.8914 to 0.8517 and the Average Contour Distance (ACD) improving from 0.8639 to 0.8263. These results demonstrate SPCL’s ability to sharpen class separation, by refining the model’s feature representations. The complete potential of SPCL becomes particularly apparent when integrated with complementary architectural components. For example, when combined with the Multi-Scale Dilated Bottleneck (baseline+MSD+SPCL configuration), SPCL enables significant performance enhancements. The F1-score and Dice coefficient rise to 87.726% and 87.680%, respectively, while the ASD and ACD metrics demonstrate further improvement, reaching 0.7689 and 0.7509. These results represent a marked advancement over the baseline performance. This combination highlights SPCL’s ability to complement multi-scale feature extraction, enabling the model to distinguish between classes with remarkable performance across varying scales.
4.5.2 Impact of the multi-scale dilated bottleneck (MSD-bottleneck)
@@
The MSD-Bottleneck module employs dilated convolutions with varying dilation rates to capture multi-scale features, thereby enabling robust handling of objects with diverse sizes and morphological characteristics, a critical capability for coronary DSA image segmentation. When integrated into the baseline architecture, the MSD-Bottleneck produces substantial performance gains, increasing recall from 85.212% to 86.530% while simultaneously reducing both the average surface distance (ASD) from 0.8914 to 0.7989 and the average contour distance (ACD) from 0.8639 to 0.8136. These quantitative improvements underscore the fundamental importance of multi-scale feature extraction for accurate segmentation of vascular structures across different spatial scales.
@@
The module’s effectiveness is further amplified when combined with the Supervised Prototypical Contrastive Loss (SPCL), as evidenced by additional reductions in segmentation error metrics to 0.7689 (ASD) and 0.7509 (ACD). This complementary interaction between MSD-Bottleneck and SPCL arises from their complementary mechanisms: while MSD-Bottleneck provides comprehensive multi-scale feature extraction through its dilated convolution architecture, SPCL concurrently optimizes the discriminative quality of these features. The combined operation of these components yields superior segmentation performance compared to their individual contributions, demonstrating that multi-scale feature capture and feature space refinement operate in a mutually reinforcing manner.
@@ 4.5.3 Role of the contextual attention fusion module (CAFM)
The CAFM Module enhances segmentation performance by selectively integrating contextual information from diverse image regions while suppressing irrelevant features. This attention mechanism substantially improves the model’s global context comprehension, a critical factor for precise coronary vessel segmentation. When implemented with the baseline architecture (baseline+CAFM), the module demonstrates measurable performance improvements: Recall increases from 85.212% to 85.990%, accompanied by rises in both F1-score (87.233% to 87.593%) and Dice coefficient (87.181% to 87.547%). Furthermore, CAFM reduces segmentation errors substantially, as evidenced by decreases in Average Surface Distance (ASD) from 0.8914 to 0.8003 and Average Contour Distance (ACD) from 0.8639 to 0.7820, confirming enhanced boundary alignment precision. However, the module’s incremental benefits become less substantial when combined with other advanced components.
In configurations combining CAFM with either the Supervised Prototypical Contrastive Loss (baseline+SPCL+CAFM) or the Multi-Scale Dilated Bottleneck (baseline+CAFM+MSD), performance gains are more modest, showing only slight improvements in F1-scores and minimal reductions in ASD and ACD metrics. This observation suggests that while CAFM independently provides significant contextual understanding enhancements, its additive value diminishes when integrated with other sophisticated modules, as their respective contributions appear to operate through complementary rather than amplifying mechanisms with respect to CAFM’s functionality.
@@ 4.5.4 Combined contributions
The complete MSA-UNet3+ with SPCL loss architecture, integrating all three key components (MSD-Bottleneck, CAFM, and SPCL), demonstrates optimal performance across all evaluation metrics. The unified model achieves a Recall of 86.775%, along with an F1-score of 87.776% and Dice coefficient of 87.733%, while simultaneously attaining the most favorable boundary alignment metrics with ASD and ACD values of 0.7587 and 0.7411 respectively. These results illustrate the effective collaborative operation of the constituent modules: the MSD-Bottleneck’s multi-scale feature extraction capability, CAFM’s contextual information integration, and SPCL’s enhanced feature discrimination collectively establish a comprehensive and precise segmentation framework. Each component addresses specific challenges in coronary vessel segmentation, and their combined operation enables superior performance in handling complex anatomical variations. Through systematic ablation analysis, we confirm that every module (MSD-Bottleneck, CAFM, and SPCL) makes substantial and distinct contributions to the overall performance of MSA-UNet3+. The complete integration of these components yields segmentation results that surpass current state-of-the-art approaches, positioning MSA-UNet3+ as a particularly effective solution for the demanding requirements of coronary DSA image analysis.
@@ 5 Conclusion
In this study, we introduced the MSA-UNet3+, a novel architecture designed to enhance coronary Digital Subtraction Angiography (DSA) image segmentation. The framework integrates a multi-scale dilated bottleneck (MSD-bottleneck) with contextual attention fusion modules (CAFM) to enable precise multi-scale feature extraction while preserving fine vessel details—critical for capturing complex vascular structures. Our technique is centered on the Supervised Prototypical Contrastive Loss (SPCL), which significantly improves the model’s ability to handle class imbalance and high intra-class variance in coronary DSA images. By integrating supervised contrastive learning with prototypical contrastive learning, SPCL effectively clusters similar features while distinguishing challenging background samples. This dual mechanism not only enhances the semantic embeddings of the encoder but also ensures that the model focuses on learning the most informative features for segmentation, particularly foreground and hard-to-classify background samples, complementing the architecture’s multi-scale capabilities. Our experimental results demonstrate that MSA-UNet3+ outperforms existing state-of-the-art methods, achieving a Dice coefficient of 87.73% and an F1-score of 87.78%. The architecture’s ability to maintain robust performance in challenging imaging conditions underscores its clinical relevance in diagnosing coronary artery diseases.
@@
However, the current framework faces limitations in severe noise, motion artifacts, or extremely low contrast, where vascular structures become indistinct. These challenging scenarios currently lead to suboptimal segmentation outcomes due to the inherent limitations of angiographic imaging. Future research directions will focus on three key improvements: First, we will investigate multimodal fusion approaches incorporating complementary imaging modalities such as intravascular ultrasound (IVUS) and optical coherence tomography (OCT) to enhance segmentation reliability in problematic cases. Second, we plan to develop optimized lightweight variants of the architecture suitable for deployment on medical imaging devices, thereby facilitating real-time clinical decision-making during interventional procedures. Third, although the current implementation specifically targets coronary DSA analysis, the underlying methodology shows considerable potential for adaptation to other medical segmentation tasks characterized by similar challenges, including class imbalance and complex anatomical backgrounds.
CRediT authorship contribution statement
Rayan Merghani Ahmed: Conceptualization, Methodology, Software , Visualization, Writing – original draft. Adnan Iltaf: Writing – review & editing. Mohamed Elmanna: Validation, Writing – review & editing Gang Zhao: Data Curation, labeling. Hongliang Li: Data Curation, labeling. Du Yue: Project Administration Bin Li: Writing – review & editing, Supervision. Shoujun Zhou: Writing – review & editing, Supervision, Funding acquisition.
Declaration of competing interest
The authors declare that they have no conflicts of interest.
Acknowledgements
@@
This work was supported by the Natural Science Foundation of Guangdong Province [No. 2023A1515010673]; in part by the Shenzhen Science and Technology Innovation Bureau key project [No. JSGG20220831110400001, No. CJGJZD20230724093303007, No. KJZD20240903101259001]; in part by the Shenzhen Medical Research Fund [No. D2404001]; in part by the Shenzhen Engineering Laboratory for Diagnosis & Treatment Key Technologies of Interventional Surgical Robots [XMHT20220104009]; and the Key Laboratory of Biomedical Imaging Science and System, CAS, for the research platform support.
Data availability